OpenStudy (destinymasha):

Solve the following system of equations. 6x - 5y = - 12 - 3x + 2y = 9 A. (7, 6) B. (6, 7) C. (- 6, - 7) D. (- 7, - 6)

OpenStudy (anonymous):

Fastest way: 6x - 5y = - 12 (1) -3x + 2y = 9 (2) Multipli the ecuation (2) by 2, you'll get -6x + 4y = 18 (3) Add ecuations 1 and 3 to cancel x 0x -y = 6 Then y=-6 Then substitute y by -6 in any equation and solve for x.

OpenStudy (destinymasha):

k

OpenStudy (destinymasha):

x = -7 ?

OpenStudy (anonymous):

Right!, good work.

OpenStudy (destinymasha):

yay thanks